# SplineSketch: Even More Accurate Quantiles with Error Guarantees

### v1

@@ B.3. Proof of the Uniform Error Guarantee
The error at threshold $\tau_{i}$ is increased either due to splitting a bucket or due to taking a union of buckets when merging sketches. We first analyze the error from the splits. For the threshold $\tau_{i}$, we consider the following sequence of tree nodes $t_{0},t_{1},t_{2},\dots$ and thresholds $\sigma_{0},\sigma_{1},\sigma_{2},\dots$: Let $\sigma_{0}:=\tau_{i}$ and let $t_{0}$ be the split-node such that the split creates $\tau_{i}$ and there is no node on the path from $t_{0}$ to the root that removes $\tau_{i}$ by a join of buckets. Note that the rank error at $\tau_{i}$ will not increase after the split represented by $t_{0}$, by the description of the algorithm. Then for $a=1,2,\dots$, we define $\sigma_{a}$ and $t_{a}$ as follows: If $\sigma_{a-1}$ is created by a split of a bucket, consider the split of a bucket $(\tau^{\prime},\tau^{\prime\prime}]$ that creates $\sigma_{a-1}=(\tau^{\prime}+\tau^{\prime\prime})/2$ during the consolidate method at $t_{a-1}$. We set $\sigma_{a}:=\tau^{\prime}$ and let $t_{a}$ be the split-node in the subtree of $t_{a-1}$ such that the split at $t_{a}$ creates $\sigma_{a}$ and $t_{a}$ is the closest node to $t_{a-1}$ with the property that on the path from $t_{a}$ to $t_{a-1}$, there is no join of buckets that removes $\sigma_{a}$. Otherwise, $\sigma_{a-1}$ is created by evenly spaced selection of initial buckets at $t_{a-1}$ (i.e., the sketches represented in the subtree of $t_{a-1}$ have no buckets, except for $t_{a-1}$), and we stop this process. Let $m$ be the index when this process stops; that is, at $t_{m}$, threshold $\sigma_{m}$ was created by the evenly spaced selection.
@@
Appendix C Reproducibility of Experimental Evaluation
Here, we provide more detailed information on running our benchmarks. The code to run the experiments, including prototypes of SplineSketch in Python and Java and code to generate the synthetic datasets, is available at https://github.com/PavelVesely/SplineSketch-experiments. Experiments were performed on an AMD EPYC 7302 (3 GHz) server with 251 GB RAM and SSD storage.
Quantile sketches.
We have used the Java prototype of SplineSketch. We have also evaluated $t$-digest [Dunning and Ertl, 2019] (v3.3, https://github.com/tdunning/t-digest), MomentSketch [Gan et al., 2018] (https://github.com/stanford-futuredata/msketch), and KLL [Karnin et al., 2016] (Java implementation by Apache DataSketches, v6.0.0, https://github.com/apache/datasketches-java). For $t$-digest, we use the default merging variant with $k_{0}$ scale function that aims at the uniform error [Dunning and Ertl, 2019]. KLL was used in the variant with the double data type, i.e., an instance of KllDoublesSketch. We note that the MomentSketch implementation only answers quantile queries, while our experiments required rank estimates; we have simulated rank queries using binary search that was done for all $m$ queries in parallel in order to decrease the number of calls to the quantile query method from $\approx m\log_{2}n$ (querying just one rank) to $\approx\log_{2}n$ calls (querying $m$ ranks), to avoid repeated computation of the distribution fitting the moments and log-moments (which is a costly operation).
The sketch size is measured in bytes when serialized on disk for storage without supporting data structures such as buffers. More precisely, for SplineSketch and $t$-digest we count 16 bytes per bucket/centroid, for MomentSketch we measured $1616k+$ 16 bytes ($k$ moments and log-moments, minimum, and maximum), and for KLL we count 8 bytes per stored item.
Datasets.
[…]
Rank queries are generated by equally spaced selection from sorted data. The queries are executed in one batch whenever possible.
Appendix D Figures with Complete Experimental Evaluation
We provide additional figures with experimental results. Figs. 6 and 7 contain the update and query times depending on the sketch size. Fig. 8 and 9 contains complete results for the average and maximum error, respectively.
Figure 6. Time per update in $\mu$s depending on sketch size on $n=10^{8}$ samples from normal distribution.
Figure 7. Time per query in $\mu$s depending on sketch size for $10^{5}$ queries on $n=10^{8}$ samples from normal distribution.
(a) HEPMASS dataset [Whiteson, 2016]
(b) Power dataset [Hebrail and Berard, 2006]
(c) Books dataset [Kipf et al., 2019, Marcus et al., 2020]
(d) Normal distribution
(e) Uniform distribution
(f) Pareto distribution
(g) Lognormal distribution
(h) Loguniform distribution
(i) Signed loguniform distribution with large exponents
(j) Normal distribution with a small parameter change after $n/2$ items
(k) Normal distribution with a large parameter change after $n/2$ items
(l) Normal distribution with $n/2$ items, followed by $n/2$ high-frequency items
Figure 8. Average rank error (log-scale) depending on the sketch size in bytes. The average is taken over $10^{5}$ evenly spaced queries.
(a) HEPMASS dataset [Whiteson, 2016]
(b) Power dataset [Hebrail and Berard, 2006]
(c) Books dataset [Kipf et al., 2019, Marcus et al., 2020]
(d) Normal distribution
(e) Uniform distribution
(f) Pareto distribution
(g) Lognormal distribution
(h) Loguniform distribution
(i) Signed loguniform distribution with large exponentslabel
(j) Normal distribution with a small parameter change after $n/2$ items
(k) Normal distribution with a large parameter change after $n/2$ items
(l) Normal distribution with $n/2$ items, followed by $n/2$ high-frequency items
Figure 9. Maximum rank error (log-scale) depending on the sketch size in bytes. The maximum is taken over $10^{5}$ evenly spaced queries.

### v2

@@ B.3. Proof of the Uniform Error Guarantee
The error at threshold $\tau_{i}$ is increased either due to splitting a bucket or due to taking a union of buckets when merging sketches. We first analyze the error from the splits. For the threshold $\tau_{i}$, we consider the following sequence of tree nodes $t_{0},t_{1},t_{2},\dots$ and thresholds $\sigma_{0},\sigma_{1},\sigma_{2},\dots$: Let $\sigma_{0}:=\tau_{i}$ and let $t_{0}$ be the split-node such that the split creates $\tau_{i}$ and there is no node on the path from $t_{0}$ to the root that removes $\tau_{i}$ by a join of buckets. Note that the rank error at $\tau_{i}$ will not increase after the split represented by $t_{0}$, by the description of the algorithm. Then for $a=1,2,\dots$, we define $\sigma_{a}$ and $t_{a}$ as follows: If $\sigma_{a-1}$ is created by a split of a bucket, consider the split of a bucket $(\tau^{\prime},\tau^{\prime\prime}]$ that creates $\sigma_{a-1}=(\tau^{\prime}+\tau^{\prime\prime})/2$ during the consolidate method at $t_{a-1}$. We set $\sigma_{a}:=\tau^{\prime}$ and let $t_{a}$ be the split-node in the subtree of $t_{a-1}$ such that the split at $t_{a}$ creates $\sigma_{a}$ and $t_{a}$ is the closest node to $t_{a-1}$ with the property that on the path from $t_{a}$ to $t_{a-1}$, there is no join of buckets that removes $\sigma_{a}$. Otherwise, $\sigma_{a-1}$ is created by evenly spaced selection of initial buckets at $t_{a-1}$ (i.e., the sketches represented in the subtree of $t_{a-1}$ have no buckets, except for $t_{a-1}$), and we stop this process. Let $m$ be the index when this process stops; that is, at $t_{m}$, threshold $\sigma_{m}$ was created by the evenly spaced selection (Algorithm 1).